A heart defibrillator being used on a patient has an time constant of 15.0 ms due to the resistance of the patient and the capacitance of the defibrillator. a. If the defibrillator has a 5.50µF capacitance, what is the resistance of the path through the patient? (You may neglect the capacitance of the patient and the resistance of the defibrillator.) b. If the initial voltage is 10.0 kV, how long does it take to decline to 120V?
Answer #1

15 x 10 -3 Aus z = RC R= RE ca) 5.5%).-6 R. 272.7.27. --4/15x143 e (b) V = Voe-t/2 120 loxio 3 t = 0.0663454 -t 15X10-3 4.422